Fred bought a new computer on sale for 30% off the original price. The store charged him 7% sales tax and his total cost was $749.00. What was the original price of the computer.
Answer by stanbon(48545)

Fred bought a new computer on sale for 30% off the original price. The store charged him 7% sales tax and his total cost was $749.00. What was the original price of the computer.
---------------
Final Price: 749
Before 7% sales tax: 749/1.07 = 700
Before 30% discount:
------------
Original Price = 700/0.7 = $1000
-------------
If you need an equation; let x be the original price
EQUATION:
749 = 1.07*0.7x
x = 749/0.749
x = 1000
==============
Cheers,
Stan H.